OpenStudy (anonymous):

solve the equation 6561^(2n)=729^(n+2)

Parth (parthkohli):

= (9^4) ^ 2n = (9 ^ 3) ^ n + 2 = 3 ^ 8n = 3 ^ 3n + 6 = 8n = 3n + 6 => 5n = 6 => n = 6/5 => n = 1.20

OpenStudy (anonymous):

n = 1/2

Parth (parthkohli):

No... it's not 1/2

OpenStudy (anonymous):

so u hav to find a common nmbr in order to wrk wth it?

Parth (parthkohli):

You can check on any calculator. n = 6/5..

Parth (parthkohli):

@wolfgirl82 Yes, get the base same, which in this case is 9, then you could make an equation of powers.

OpenStudy (anonymous):